# Simulation of time crystal behavior for a few boson chiral soliton model in a ring

### 09 Jan 2024 v1

@@ I Introduction
Since Wilczek’s seminal paper initiated the field of quantum time crystals in 2012 Wil12 , it has attracted intense interest, both fundamental and applied. His model comprised $N$ bosons with attractive interactions in a ring that was pierced by a flux tube, the key idea being that this model could yield a ground state composed of a soliton that was rotating around the ring. Although this model was subsequently shown to always yield a non-rotating ground state, the time crystal idea had taken root. Now there is active research in both continuous and discrete time crystals and for a variety of platforms including optics, ultracold atoms, metamaterials, spin systems, and both closed and open systems. For recent reviews of the area see SacZak18 ; Sacha20 ; ZalLukMon23 .
@@
The remainder of this paper is organized as follows: Section II describes the few boson version of our chiral soliton model along with a description of the parameters involved. The case of $N=2$ particles is covered in Sec. III and shows the ideas involved in a transparent manner. Section IV contains the main results and covers the three particle case that demonstrates that our few boson chiral soliton model can display time crystal behavior in which a two-boson soliton state can execute several revolutions of the ring. Summary and conclusions are given in Sec. V.
@@ II.1 N-particle Schrödinger equation
We consider a system of $N$ scalar bosons of mass $m$ located circumferentially in a ring of radius $R$. The Schrödinger equation for the N-particle wavefunction $\psi(\theta_{1},\theta_{2},\ldots\theta_{N},t)$ is in scaled units Jackiw
@@ III.2 Two-particle ground state
with $\varepsilon$ the scaled energy of the two-particle system, and $p$ an even integer that is the winding number associated with the center-of-mass OAM of the two particles. Substituting this ansatz into Eq. (7) yields
@@ III.3 Time crystal dynamics from measurement
For time crystal behavior what is required is that the center of the initial localized state basically translates in time and persists in shape for several rotations around the ring. The speed and direction of the rotation are intimately tied to the chiral gauge potential proportional to $\kappa$. Figure 4 shows an example of time crystal behavior using $n=1$ and $\Delta\theta\sim 2$, which means a much weaker measurement than in Figure 3(b). In this case the uncertainty is greater than the expected width $w\sim 1$ for a two-particle localized state, so the measurement is expected to be less disruptive. This figure shows that the initial localized solution remains largely intact while it rotates, with a velocity $v\sim-1$, and it persists over several rotations around the ring. The initial position of the centroid of the localized solution is set by the outcome of the initial measurement, $\theta^{(0)}=0$ in Fig. 4(b). In addition, if one selects a value for $\theta_{1}\neq 0$ in Fig. 4 the resulting time variation of the probability density is then largely same for any other value of $\theta_{1}$, just displaced in time. This example therefore exhibits time crystal behavior to a good approximation. This may seem counter-intuitive at first, since one would expect a localized initial state to expand and eventually interfere with itself in the ring giving rise to complicated dynamics. However, it is possible to construct special solutions to the single particle Schrödinger equation for a ring, that are localised non-dispersive solutions with a non-zero angular momentum, i.e. they rotate. Such solutions have been found by Qin Qin and are of the form
$\Phi(\theta,t)=C_{0}\cos(m\theta-nt)e^{i(m\theta-{1\over 2}(m^{2}+n^{2})t)},$ (14)
where $m$ and $n$ are integers. From Eq. (14) we notice that $\Phi(\theta,t)$ is a travelling localized solution with velocity $v=n/m$. For $m=-n=1$ we find $v=-1$, in qualitative agreement with the results in Fig. 4. In addition, because of the linearity of the single-particle Schröinger equation, we can always construct superpositions of different states such that we have a non-zero background. This is indeed the type of solution we generally find after measuring one particle with sufficiently large uncertainty. We therefore conclude that when extracting information about the two-particle ground state by measuring the position of one particle with sufficiently large uncertainty, the ground state is projected onto a single particle wave function which is localised, shape preserving, and rotating in the ring.
@@ IV Three-particle case
Although the two-particle case provides some insight it does not speak to the bigger issue of whether the spontaneously formed and rotating solution after the measurement can persist even for a multi-particle state, so the rotating solution may be viewed even approximately as a rotating soliton. For this reason we next consider the three-particle case so that two particles are left on the ring after the measurement, with the possibility of exciting the two-particle analogue of a soliton.
@@ IV.1 Three-particle Schrödinger equation
$i{\partial\Psi\over\partial t}=-\left(2{\partial^{2}\over\partial x^{2}}+6{\partial^{2}\over\partial y^{2}}+{1\over 3}{\partial^{2}\over\partial s^{2}}\right)\Psi-2i\kappa\rho(x,y){\partial\Psi\over\partial s}+3[\kappa\rho(x,y)]^{2}\Psi+{g\over 2}\rho(x,y)\Psi.$ (21)
@@ IV.2 Three-particle ground state
where $p$ is the winding number associated with the center-of-mass OAM of the three particles: As a result $p$ should be an integer multiple of three. Substituting the above ansatz into Eq. (21) yields
$\varepsilon\varphi(x,y)=-\left(2{\partial^{2}\over\partial x^{2}}+6{\partial^{2}\over\partial y^{2}}\right)\varphi+3\left[{p\over 3}+\kappa\rho(x,y)\right]^{2}\varphi+{g\over 2}\rho(x,y)\varphi,$ (23)
@@
We have solved Eq. (23) numerically for a range of parameters using the imaginary time method LehToiElo2007 , in particular to obtain the ground state $\varphi(x,y)$ and associated scaled energy $\varepsilon$. Figure 5 shows numerical results for parameters $\ell=3,g=-2,\kappa=0.2,q=50$. Specifically, Fig. 5(a) shows the scaled energy versus $p$ and reveals that the ground state occurs for $p=\ell=3$, and (b) shows a line plot of the ground state profile $|\varphi(x,y)|^{2}$ versus $x$ and $y$ for $p=3$.
Figure 5: (a) Scaled energy $\varepsilon$ versus COM winding number $p$, and (b) ground state profile $|\varphi(x,y)|^{2}$ for $p=\ell=3$. Parameters are $\ell=3,g=-2,\kappa=0.2,q=50$.
@@ IV.3 Time crystal dynamics from measurement
In their work SKS argued that a precise position measurement of one particle could lead initially to spontaneous formation of a localized solution SKS_EPL . We assume that the position of particle $3$ is measured yielding a value $\theta_{3}^{(0)}$, but allow for the measurement to be imprecise, so that the wave function for the remaining particles may be written as
@@
For time crystal behavior what is required is that the center of the initial localized state basically translates in time and persists in shape for several rotations around the ring. The speed and direction of the rotation are intimately tied to the chiral gauge potential proportional to $\kappa$. Figure 7 shows an example of time crystal behavior using $n=1$ and $\Delta\theta\sim 2$, which means a much weaker measurement than in Figure 6(b): In this case the uncertainty is greater than the expected width $w\sim 1$ for a three-particle localized state, so the measurement is expected to be less disruptive. This figure shows that the initial localized solution remains largely intact while it rotates, and it persists over several rotations around the ring. In addition, if one selects a value for $\theta_{1}$ in Fig. 7 the resulting time variation of the probability density is then largely same for any other value of $\theta_{1}$, just displaced in time. This example therefore exhibits time crystal behavior to a good approximation.
@@
Figure 6: (a) Reconstructed ground state probability density $|\Psi(\theta_{1},\theta_{2},\theta_{3}=0)|^{2}$, and (b) simulation for a measurement with $n=50$ that localizes particle $3$ with an uncertainty $\Delta\theta\sim 0.2$. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it does not reflect time crystal behavior. Parameters are the same as as in Fig. 5 along with $g=-2$.
Figure 7: Simulation for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2$ that couples the remaining two particles after the measurement of particle $3$ to the two-particle limit of the chiral soliton. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it is quite periodic, consistent with time crystal behavior over several periods. Parameters are the same as as in Fig. 5 along with $g=-2$.
@@ V Summary and conclusions
In 2019 we proposed a mean-field chiral soliton model for a quantum time crystal OW19 , and this proposal met with some criticism in the literature SKS_EPL ,SKS_comment -OW_lack . The goal of the present paper was to examine a few boson $(N=2,3)$ limit of our previous chiral soliton model to numerically assess whether time crystal behavior is possible in the context of the type of calculation in Ref. SKS_EPL . In particular, we initialize the system in the $N$-boson ground state, perform a position measurement of one particle, and see if spontaneous formation of the resulting $(N-1)$-boson localized state (soliton for $N=2$) can persist over several revolutions of the ring as required for time crystal behavior. We find that for an imprecise or weak position measurements time crystal behavior is possible in our chiral soliton model, whereas for a precise measurement quantum fluctuations cause the soliton to decay in accordance with SKS_EPL .
@@
Clearly the small particle numbers and parameters used here are not compatible with experiments, but are rather intended to show that time crystal behavior is not ruled out in our model. We remark that the same behavior could be seen for other combinations of parameters as long as a weak measurement is employed, with $\Delta\theta>w$. Restricting to a few bosons allowed for numerical simulations to be performed for our chiral soliton model without undue approximations, but our conclusions may still be of relevance in the more general case with large $N$. According to the Wilczek model analyzed in Ref. SKS_EPL the soliton lifetime $t_{c}$ is expected to scale with $N$. Assuming that this scaling also applies to our chiral model, if $N=3$ can show time crystal behavior then it should be expected for larger $N$ as the lifetime increases. Therefore we argue that the chiral model can realize a quantum time crystal when weak position measurements are incorporated into the scheme.
Figure 8: Simulation for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2$ that couples the remaining particle after the measurement of particle $2$ to a linear non-dispersive wave. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it is quite periodic, consistent with time crystal behavior over several periods. Parameters are the same as in Fig. 5 along with $g=-2$.
@@
Finally, we discuss the issue of whether our simulations point to genuine time crystal behavior. This would require that the periodic motion seen in Fig. 7 would persist for all times for $N=3$, and we cannot be assured that this is the case based on numerical solutions. What we can say for sure is the time crystal behavior can persist over several revolutions of the ring if imprecise measurements are allowed for. Specifically, the first position measurement sets the time crystal in motion and establishes a time reference by setting the initial position of the $N=2$ soliton peak, but further measurements will be required to monitor if the time crystal continues to exhibit periodic motion. Indeed, for $N=3$ we can go one step further and ask whether time crystal behavior would persist if we performed a second imprecise measurement, of say particle $2$, acting on the two-particle state in Fig. 7. Figure 8 shows a simulation for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2$, along with $\theta_{2}^{(0)}$ chosen consistent with the probability density for such a measurement between $[-\pi,\pi]$, that couples the remaining particle at $t=20$ to a linear non-dispersing state after the measurement of particle $2$. Then picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it is quite periodic, consistent with time crystal behavior over several periods. For our $N=3$ case this validates that the time crystal behavior can survive more than one measurement. In future publications we plan to extend this initial numerical study to larger particle numbers, both numerically and analytically, and thereby move the project towards experimentally accessible parameter ranges.
@@ VI Acknowledgements
We acknowledge helpful discussions with Erika Anderson, Stewart Lang, Joel Priestley, Gerard Valentí-Rojas, and Brian Anderson.

### 13 Mar 2024 v3

@@ I Introduction
Since Wilczek’s seminal paper initiated the field of quantum time crystals in 2012 Wil12 , it has attracted intense interest, both fundamental and applied. His model comprised $N$ bosons with attractive interactions in a ring that was pierced by a flux tube, the key idea being that this model could yield a ground state composed of a soliton that was rotating around the ring. Although this model was subsequently shown to always yield a non-rotating ground state Bruno , the time crystal idea had taken root. Now there is active research in both continuous and discrete time crystals and for a variety of platforms including optics, ultracold atoms, metamaterials, spin systems, and both closed and open systems. For recent reviews of the area see SacZak18 ; Sacha20 ; ZalLukMon23 .
@@
The remainder of this paper is organized as follows: Section II describes the few boson version of our chiral soliton model along with a description of the parameters involved. The case of $N=2$ particles is covered in Sec. III and shows the ideas involved in a transparent manner. Section IV contains the main results and covers the three particle case that demonstrates that our few boson chiral soliton model can display time crystal behavior in which a two-boson soliton state Jackiw can execute several revolutions of the ring. Summary and conclusions are given in Sec. V.
@@ II.1 N-particle Schrödinger equation
We consider a system of $N$ scalar bosons of mass $m$ located circumferentially in a ring of radius $R$. The Schrödinger equation for the N-particle wavefunction $\psi(\theta_{1},\theta_{2},\ldots\theta_{N},t)$ is in scaled units PriValWri23
@@ III.2 Two-particle ground state
with $\varepsilon$ the scaled energy of the two-particle system, and $p$ an integer that is the winding number associated with the center-of-mass OAM of the two particles. Substituting this ansatz into Eq. (7) yields
@@ III.3 Time crystal dynamics from measurement
For time crystal behavior what is required is that the center-of-mass of the initial localized state translates in time while maintaining its shape for several transits of the ring. An estimate of the magnitude of the rotation rate may be obtained as follows: Assuming some portion of the center-of-mass OAM pPlanck-constant-over-2-$\hbar p$ of the ground state is transferred to the localized state via the weak measurement, we find that $mNR|{ds\over dt}|\leq\hbar|p|$, with ${ds\over dt}$ the velocity of the center-of-mass of the localized state, and the rotation rate obeys $|\Omega|={1\over R}|{ds\over dt}|\leq{\hbar|p|\over mNR^{2}}$. In our scaling the unit for time is $t_{s}={\hbar\over 2mR^{2}}$, so the scaled rotation rate obeys $|v|=|\Omega|t_{s}\leq{2|p|\over N}$, or $|v|\leq 2$ for the current example. Figure 4 shows an example of time crystal behavior using $n=1$ and $\Delta\theta\sim 2$, which means a much weaker measurement than in Figure 3(b). In this case the uncertainty is greater than the expected width $w\sim 1$ for a two-particle localized state, so the measurement is expected to be less disruptive. This figure shows that the initial localized solution remains largely intact while it rotates with a velocity $v\sim 1$ Neq2 , and it persists over several rotations around the ring. The initial position of the centroid of the localized solution is set by the outcome of the initial measurement, $\theta^{(0)}=0$ in Fig. 4(b). In addition, if one selects a value for $\theta_{1}\neq 0$ in Fig. 4 the resulting time variation of the probability density is then largely same for any other value of $\theta_{1}$, just displaced in time. This example therefore exhibits time crystal behavior.
The above argument may appear counter-intuitive since one would expect a localized initial state to expand and produce interference around the ring, giving rise to complicated dynamics. However, it is possible to construct solutions to the single particle Schrödinger equation for a ring that are localised, non-dispersive, and rotating that have non-zero angular momentum Qin . Adapted to the solution of Eq. (13) these solutions take the form
$\psi(\theta_{1},t)={\cal N}_{0}\cdot\cos(m(\theta_{1}-vt))e^{i\ell\theta_{1}/2+iv\theta_{1}/2-i(m^{2}+v^{2}/4)t},$ (14)
where ${\cal N}_{0}$ is a normalization constant, and $m$ and $v$ are solution parameters. From Eq. (14) we notice that $\psi(\theta_{1},t)$ is a travelling localized solution with velocity $v$, and $2m$ is the number of wave function nodes around the ring. For the simulation in Fig. 4 for $\ell=2$ there is one node around the ring, so $m=1/2$, and the velocity is $v=1$. Moreover, one can verify that the corresponding initial wave function
$\psi(\theta_{1},t=0)={\cal N}_{0}\cdot\cos\left({\theta_{1}\over 2}\right)e^{3i\theta_{1}/2},$ (15)
is a physically allowed single-valued wave function. We have verified numerically that when the weak measurement is performed on the two-particle ground state using Eq. (11), the initialized wave function is well approximated near the origin by the above wave function $\psi(\theta_{1},t=0)$. The physical relevance of this is that, given our choice of measurement distribution in Eq. (12) with $n=1$ and $\theta^{(0)}_{2}=0$, the weak measurement dominantly excites a travelling wave solution of the ring that does not spread. In addition, because of the linearity of the single-particle Schrödinger equation, we can always construct superpositions of states such that we have a non-zero background as appears in the simulations.
@@ IV Three-particle case
Although the two-particle case provides some insight it does not speak to the bigger issue of whether the spontaneously formed and rotating solution after the measurement can persist even for a multi-particle state, so the rotating solution may be viewed even approximately as a rotating soliton. For this reason we next consider the three-particle case so that two particles are left on the ring after the measurement, with the possibility of exciting the two-particle analogue of a soliton Jackiw .
@@ IV.1 Three-particle Schrödinger equation
$i{\partial\Psi\over\partial t}=-\left(2{\partial^{2}\over\partial x^{2}}+6{\partial^{2}\over\partial y^{2}}+{1\over 3}{\partial^{2}\over\partial s^{2}}\right)\Psi+2i\kappa\rho(x,y){\partial\Psi\over\partial s}+3\left[{\ell\over 2}+\kappa\rho(x,y)\right]^{2}\Psi+{g\over 2}\rho(x,y)\Psi.$ (22)
@@ IV.2 Three-particle ground state
where $p$ is the integer valued winding number associated with the center-of-mass OAM of the three particles. Substituting the above ansatz into Eq. (22) yields
$\varepsilon\varphi(x,y)=-\left(2{\partial^{2}\over\partial x^{2}}+6{\partial^{2}\over\partial y^{2}}\right)\varphi+3\left[{p\over 3}-{\ell\over 2}-\kappa\rho(x,y)\right]^{2}\varphi+{g\over 2}\rho(x,y)\varphi,$ (24)
@@
We have solved Eq. (24) numerically for a range of parameters using the imaginary time method LehToiElo2007 , in particular to obtain the ground state $\varphi(x,y)$ and associated scaled energy $\varepsilon$. Figure 5 shows numerical results for parameters $\ell=2,g=-0.8,\kappa=0.1,q=50$. Specifically, Fig. 5(a) shows the scaled energy versus $p$ and reveals that the ground state occurs when ${p\over N}={\ell\over 2}$, or $p=3$ for this example, and (b) shows a line plot of the ground state profile $|\varphi(x,y)|^{2}$ versus $x$ and $y$ for $p=3$.
Figure 5: (a) Scaled energy $\varepsilon$ versus COM winding number $p$, and (b) ground state profile $|\varphi(x,y)|^{2}$ for $p=3$. Parameters are $\ell=2,g=-0.8,\kappa=0.1,q=50$.
@@ IV.3 Time crystal dynamics from measurement
We next extend the measurement model to the case with three initial particles. In particular, we assume that the position of particle $3$ is measured yielding a value $\theta_{3}^{(0)}$, but allow for the measurement to be imprecise, so that the wave function for the remaining particles may be written as
@@
Figure 6: (a) Reconstructed ground state probability density $|\Psi(\theta_{1},\theta_{2},\theta_{3}=0)|^{2}$, and (b) simulation for a measurement with $n=50$ that localizes particle $3$ with an uncertainty $\Delta\theta\sim 0.2$. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it does not reflect time crystal behavior. Parameters are the same as as in Fig. 5.
For time crystal behavior what is required is that the center of the initial localized state basically translates in time and persists in shape for several rotations around the ring. We may use our previous estimate for the magnitude of the scaled rotation rate $|v|\leq{2|p|\over N}$, or $|v|\leq 2$ for the current example. Figure 7 shows an example of time crystal behavior using $n=1$ and $\Delta\theta\sim 2$, which means a much weaker measurement than in Figure 6(b): In this case the uncertainty is greater than the expected width $w\sim 1$ for a three-particle localized state, so the measurement is expected to be less disruptive. This figure shows that the initial localized solution remains largely intact while it rotates with velocity $v\sim-1$ Neq3 , and it persists over several rotations around the ring. In addition, if one selects a value for $\theta_{1}$ in Fig. 7 the resulting time variation of the probability density is then largely same for any other value of $\theta_{1}$, just displaced in time. This example therefore exhibits time crystal behavior to a good approximation.
@@
Figure 7: Simulation for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2$ that couples the remaining two particles after the measurement of particle $3$ to the two-particle limit of the chiral soliton. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it is quite periodic, consistent with time crystal behavior over several periods. Parameters are the same as as in Fig. 5.
Figure 8: (a) Scaled energy $\varepsilon$ versus COM winding number $p$, and (b) ground state profile $|\varphi(x,y)|^{2}$ for $p=6$. Parameters are $\ell=4,g=-1,\kappa=0.1,q=50$.
@@ V Summary and conclusions
In 2019 we proposed a mean-field chiral soliton model for a quantum time crystal OW19 , and this proposal met with some criticism in the literature SKS_EPL ,SKS_comment -OW_lack . The goal of the present paper was to examine a few boson $(N=2,3)$ limit of our previous chiral soliton model to numerically assess whether time crystal behavior is possible in the context of the type of calculation in Ref. SKS_EPL . In particular, we initialize the system in the $N$-boson ground state, perform a position measurement of one particle, and see if spontaneous formation of the resulting $(N-1)$-boson localized state (soliton for $N=3$) can persist over several revolutions of the ring as required for time crystal behavior. We find that for an imprecise or weak position measurements time crystal behavior is possible in our chiral soliton model, whereas for a precise measurement quantum fluctuations cause the soliton to decay in accordance with SKS_EPL .
One feature from the simulations is that in the presence of chirality the quantum ground state of the system can carry orbital angular momentum pPlanck-constant-over-2-$\hbar p$, with $p$ the center-of-mass winding number which can be non-zero for the state of lowest energy. This is illustrated in Fig. 2(a) for $N=2$ with $p=2$ for minimum energy, and Fig. 5(a) for $N=3$ with $p=3$, both for $\ell=2$. More generally, we find that for minimum energy ${p\over N}={\ell\over 2}$, and this is illustrated in Fig. 8(a) for $\ell=4$, giving $p=6$ for minimum energy. The point is that the presence of chirality and non-zero orbital angular momentum appear to be key requirements for time-crystal behavior to be a possibility in our model, the imprecise measurement facilitating transfer of the orbital angular momentum to the center-of-mass momentum of the localized soliton. Estimates for the rotation rates were given but details of this process remain to be elucidated.
@@
Clearly the small particle numbers and parameters used here are not compatible with experiments, but are rather intended to show that time crystal behavior is not ruled out in our model. We remark that the same behavior could be seen for other combinations of parameters as long as a weak measurement is employed, with $\Delta\theta>w$. Restricting to a few bosons allowed for numerical simulations to be performed for our chiral soliton model without undue approximations, but our conclusions may still be of relevance in the more general case with large $N$. According to the Wilczek model analyzed in Ref. SKS_EPL the soliton lifetime $t_{c}$ is expected to scale with $N$. If this scaling also applies to our chiral model, then if $N=3$ can show time crystal behavior it should be expected for larger $N$ as the lifetime increases.
@@
Finally, there is the issue of whether our simulations point to genuine time crystal behavior. This would require that the periodic motion seen in Fig. 7 would persist for all times for $N=3$, and we cannot be assured that this is the case based on numerical solutions. What we can say for sure is the time crystal behavior can persist over several revolutions of the ring if imprecise measurements are allowed for. In future publications we plan to extend this initial numerical study to larger particle numbers, both numerically and analytically, and thereby move the project towards experimentally accessible parameter ranges.
@@ VI Acknowledgements
We acknowledge helpful discussions with Brian Anderson, Erika Andersson, Stewart Lang, Joel Priestley, and Gerard Valentí-Rojas.